Q. Choose the correct answer.
Q. The equation u_xx 〖-4u〗_xy=0 is
)parabolic
b)hyperbolic
c)Elliptic
d) None of above
1
Expert's answer
2019-07-12T12:10:40-0400
"u_{xx} - 4u_{xy} =0"

The equation

"A u_{xx}+2 B u_{xy} + C u_{yy} + D u_{x} + E u_{y} + F =0"

is elliptic if

"AC-B^2 >0"

hyperbolic if

"AC - B^2 <0"

and parabolic if

"AC - B^2 =0"

We have A = 1, B = -2, C =0, hence

"1 \times 0 - (-2)^2 = -4<0"

i.e. the answer is b) hyperbolic